OpenStudy (anonymous):

Help Please ! What polynomial identity should be used to prove that 133 = 125 + 8? Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes

OpenStudy (solomonzelman):

My identity, "Common sense".

OpenStudy (anonymous):

I Don't Get How To Figure It Out . . .

OpenStudy (cwrw238):

its simple addition of course bu you can use the sum of cubes identity

OpenStudy (anonymous):

i know that 125+8= 133. however, i do not know which polynomial identity it is. so i get the addition part, im just confused about the polynomial identity's

OpenStudy (cwrw238):

a^3 + b^3 = ( a + b)(a^2 - ab + b^2)

OpenStudy (isaiah.feynman):

5^3+2^3=133

OpenStudy (anonymous):

Ohh Okayy , i get it now. Thanks Guys